i was tought this but i forgot sooo: how many grams of salt must be added to 30 grams of a 20% salt solution to obtain a 25% salt solution?
Answer by stanbon(48558)

how many grams of salt must be added to 30 grams of a 20% salt solution to obtain a 25% salt solution?
----------------------------------------------
Equation:
salt + salt = salt
0.20*30 + x = 0.25(30+x)
Multiply thru by 100 to get:
600 + 100x = 25*30 + 25x
75x = 150
x = 2 grams (amt. of salt that must be added to the mixture)
================================================================
Cheers,
Stan H.